Consider the reaction of copper (II) sulfate with sodium hydroxide.

a) Write the balanced molecular equation for this reaction.

b) Write the balanced complete ionic equation for this reaction.

c) Write the balanced net ionic equation for this reaction.

Use proper subscript and superscript formatting, and include states.
1
Expert's answer
2020-04-18T07:06:38-0400

a). CuSO4(s) +2NaOH(aq) = Cu(OH)2(aq) +Na2SO4(aq)

b).CuSO4 (s) + 2Na+ (aq)+ 2OH-(aq) = Cu2+(aq) + 2OH-(aq) +Na+(aq) + SO42-(aq)

c). CuSO4(s) = Cu2+(aq) + SO42-(aq)